# Spatial structure facilitates evolutionary rescue by cost-free drug resistance

### v1

@@ Results
Second, we consider a star structure, comprising a central deme connected to $D-1$ leaves [marrec_toward_2021]. All leaves are assumed to be equivalent. Denoting by $\gamma_{I}$ (resp. $\gamma_{O}$) the migration from a leaf to the center (resp. from the center to a leaf), we defined migration asymmetry as $\alpha=\gamma_{I}/\gamma_{O}$. We compared the star with a given outgoing migration rate from a leaf to the clique with the same total outgoing migration rate from a deme (i.e. $\gamma_{0}=(D-1)\gamma$). We observe in Figure S5 that the star gives the same survival probability to treatment as the clique and the lattice if $t_{add}<\langle t_{af\,F}\rangle$, i.e., in cases where usually not more than one deme has fixed resistance before drug addition. For larger values of drug addition time $t_{add}$, our numerical results for the star depend on $\alpha$ and significantly differ from those obtained for the clique and the lattice, although the difference remains minor. We interpret this minor difference as arising from wild-type individuals possibly re-invading demes where mutants have fixed, which is structure-dependent. See SI Section 4 for more details. Overall, our main results are robust to changing the graph on which demes are placed.
@@ Discussion
Here, we showed that spatial structure can facilitate the survival of a bacterial population to antibiotic treatment, starting from a sensitive inoculum. Indeed, the bacterial population can be rescued if antibiotic resistant mutants are present when drug is added. While the emergence of resistant bacteria by random mutations only depends on total population size and not on spatial structure, their fate can be affected by spatial structure. In the relevant case where the mutation that provides resistance is neutral or effectively neutral, its probability of fixation is increased in smaller populations. This leads to local fixation of resistant mutants in demes, which then constitute refugia of resistance mutants and allow the population to survive when drug is added. Because of this, spatial structure facilitates evolutionary rescue by cost-free drug resistance. The survival probability of the population increases when the migration rate between demes is decreased, and when the number of demes is increased. Once the population is rescued by resistance, migrations allow resistant mutants to spread in all demes. While our main results were obtained starting from a sensitive inoculum, we showed that they extend to the case where there are resistant mutants in the inoculum. While we considered a minimal model where all demes are equivalent and connected to each other with identical migration rates, bacterial populations can have diverse spatial structures, with migrations that may be more local or asymmetric. We showed that our main conclusions still hold for a lattice of demes, modeling more localized migrations, and for demes located on a star graph with asymmetric migrations.
@@
Because local fixation of neutral resistant mutants is instrumental to our effect, it involves rather long timescales. The average appearance time $\langle t_{af\,F}\rangle=1/(D\mu g)$ of a locally successful mutant in the fastest deme is of particular relevance. As it scales as the inverse mutation probability per individual and per generation $1/\mu$, it can be quite long, even considering that there are often several different mutational targets that give rise to resistance, leading to a larger effective $\mu$. However, this timescale is inversely proportional to the number $D$ of demes, meaning that it can potentially become arbitrarily small when increasing population subdivision. With a mutation probability per nucleotide and per generation of $10^{-10}$ [lee_rate_2012], assuming 10 possible mutations for the development of resistance to a given drug, gives $\mu=10^{-9}$. Then, for bacteria dividing once per hour, we find $\langle t_{af\,F}\rangle=10^{8}$ h for $D=10$, but $\langle t_{af\,F}\rangle=10^{2}$ h for $D=10^{7}$. Such large numbers of demes are realistic for instance in the case of intestinal glands [Nguyen10]. Note also that larger mutation rates, e.g. in mutators, can be relevant to the evolution of resistance [Gifford23]. While we mainly focused on neutral mutants giving resistance, our findings extend to effectively neutral mutants, and to deleterious mutants. However, in the latter case, the timescales involved would be even longer.
Here, we modeled the action of the antibiotic as preventing any division of sensitive bacteria. In other words, we considered a perfect biostatic drug. However, our main findings generalize to biocidal drugs that increase the death rate of sensitive bacteria, or to drugs which combine both modes of actions. Indeed, what happens before the addition of drug, in particular the local fixation of resistant mutants, is not impacted by the mode of action of the drug. Thus, the effect of spatial structure we evidenced here holds independently of this. The difference between the biostatic and the biocidal case lies in the decay of the sensitive bacteria once drug is added. In the biocidal case, sensitive microbes may still divide during this phase, and new resistant mutants may then appear [marrec_resist_2020, czuppon_stochastic_2023]. This is the case both in well-mixed and in structured populations, with only minor differences between them, due to the stochasticity of extinction. Thus, our main findings are robust to the mode of action of the drug. Beyond modes of action, it would be interesting to study the impact of spatial structure on multi-step drug resistance evolution [Igler21].
@@ 1.1 Appearance of a mutant that fixes in a well-mixed population
In our model, mutations from sensitive to resistant individuals happen at birth at per capita rate $\mu$. Thus, since at equilibrium $f_{S}\left(1-N/K\right)=g$, the average time it takes for a mutant to appear in a population of $N$ wild-types is:
$t_{app}=\frac{1}{N\mu g}.$ (S1)
The fixation probability of neutral mutants in a population of size $N$ is:
$p_{fix}=\frac{1}{N}.$ (S2)
Thus, the average time of appearance of a mutant that fixes in a well-mixed population is given by:
$\langle t_{af\,W}\rangle=t_{app}\times\frac{1}{p_{fix}}=\frac{1}{\mu g}.$ (S3)
@@ 1.2 Appearance of the slowest mutant that fixes in a structured population
Let us now focus on the appearance of a mutant that fixes in the slowest deme in a population composed of $D$ independent demes, in the absence of migrations. By slowest deme, we mean the deme in which a locally successful mutant (i.e. a mutant destined to fix locally in its deme of origin) takes the longest time to appear.
For a mutant to fix in the slowest deme, mutants must have fixed in all others. Let us denote by $t_{af\,S}$ (resp. $t_{af\,d}$) the time of appearance of a mutant in the slowest deme (resp. in any deme). The probability $P(t_{af\,S}\leq t)$ that $t_{af\,S}$ is smaller or equal than $t$ reads for any $t$:
@@
The appearance of a locally successful mutant is a Poisson process with rate $\mu g$ (corresponding to the inverse of Equation (S3)). Thus, the time $t_{af\,d}$ is exponentially distributed with rate $\mu g$ and probability density
$p_{d}(t_{af\,d}=t)=\mu ge^{-\mu gt}\,,$ (S5)
and we have $P(t_{af\,d}\leq t)=1-e^{-\mu gt}$. This allows us to express the probability density $p_{S}(t)$ of appearance of a mutant in the slowest deme:
$\begin{split}p_{S}(t)dt&=dp(t_{af\,S}\in[t,t+dt])=P(t_{af\,S}\leq t+dt)-P(t_{af\,S}\leq t)\\ &=\frac{dP(t_{af\,S}\leq t)}{dt}dt=D\left[P(t_{af\,d}\leq t)\right]^{D-1}\,\frac{dP(t_{af\,d}\leq t)}{dt}dt\\ &=D\left[1-e^{-\mu gt}\right]^{D-1}\mu ge^{-\mu gt}dt\,.\end{split}$ (S6)
@@ 1.3 Appearance of the fastest mutant that fixes in a structured population
$\begin{split}p_{F}(t)dt&=dp(t_{af\,F}\in[t,t+dt])=-\frac{dP(t_{af\,F}>t)}{dt}dt\\ &=-D[P(t_{af\,d}>t)]^{D-1}\frac{dP(t_{af\,d}>t)}{dt}dt\\ &=-D[1-P(t_{af\,d}\leq t)]^{D-1}\left[-\frac{dP(t_{af\,d}\leq t)}{dt}\right]dt\\ &=D\mu ge^{-\mu gtD}dt\,.\end{split}$ (S9)
Note that $p_{F}(t)$ is positive for all $t$ and normalized, as expected.
@@ 1.4 Survival probability of the population with $\gamma=0$
We checked that Equation (S11) was satisfied in our numerical simulations, within errors given by the standard error of the proportion.
1.5 Correction for stochastic extinction in the mutant spreading time
We derive here the correction accounting for the stochastic extinction of mutants, which was used in the main text to obtain the mutant colonization time in Equation (3), namely $C=(1-g/f_{R})^{-1}$. Let us consider one resistant bacterium with replication rate $f_{R}$ and death rate $g$. Let us denote by $e$. the extinction probability of its lineage. This resistant individual can die without dividing with probability $p_{D}=g/(f_{R}+g)$, giving an extinction probability is $e_{D}=1$. It can also divide before dying with probability $p_{R}=f_{R}/(f_{R}+g)$, leading to the extinction probability $e_{R}=e^{2}$. We can then write an equation for the extinction probability $e$ of the lineage of the single resistant bacterium:
$e=p_{D}\,e_{D}+p_{R}\,e_{R}=\frac{g}{f_{R}+g}+e^{2}\frac{f_{R}}{f_{R}+g}.$ (S14)
The solutions to this equation are $e=1$ or $e=g/f_{R}$ if $g<f_{R}$. Then, the survival probability reads: $1-g/f_{R}$ if $g<f_{R}$. The correction for stochastic extinction in the mutant colonization time (Equation (3)) is then a multiplicative factor $C=(1-g/f_{R})^{-1}$.
Let us consider a deme where R mutants have not fixed, in the absence of drug. Let us determine the probability $p_{pres}$ of presence of mutants that arise through mutations upon division of sensitive individuals, and whose lineage is destined to go extinct. We follow the method exposed in Ref. [marrec_resist_2020], which is based on the transition rate matrix of the process.
We distinguish each case where exactly $i$ mutants are present. In each case, the contribution to $p_{pres}$ is obtained by multiplying the probability $p_{R}$ that at least one R mutant is present, by the conditional probability $p_{R}^{c}(i)$ of having $i$ R mutants, provided that a lineage of mutants exists. We have $p_{R}=\tau_{R}^{d}\,N\mu g$, where $\tau_{R}^{d}$ is the the average lifetime of a resistant lineage destined to extinction, and $N\mu g$ the total mutation rate in the deme, with $N$ the steady-state deme size, $g$ the death rate, $\mu$ the mutation rate. Meanwhile, $p_{R}^{c}(i)=\tau_{R,i}^{d}/\tau_{R}^{d}$, with $\tau_{R,i}^{d}$ the average time the lineage spends in a state with $i$ mutants, calculated as $\tau_{R,i}^{d}=-\pi_{i}/\pi_{1}(\mathbf{\tilde{R}}^{-1})_{i1}$. Here, $\pi_{i}$ is the probability that resistant mutants go extinct, starting from $i$ of them, while $\mathbf{\tilde{R}}$ is the reduced transition rate matrix for mutants (i.e., the transition rate matrix where rows and columns corresponding to absorbing states are eliminated [marrec_resist_2020]).
Summing the term $p_{R}\,p_{R}^{c}(i)$ over all possible $i$, we obtain the probability $p_{pres}$ that mutants from a lineage destined to go extinct are present in the deme:
$p_{pres}=-\sum_{i=1}^{N}Ng\mu\,\frac{\pi_{i}}{\pi_{1}}(\tilde{\mathbf{R}}^{-1})_{i1}\,.$ (S15)
As discussed in the main text, when drug is added to the system, the population can survive if mutants are present when the drug is added. For survival, it is also necessary that the mutant lineages do not go stochastically extinct. Thus, combining $p_{pres}$ with the stochastic extinction probability (see Subsection 1.5) yields an estimate of the survival probability $p_{surv}$ assuming that mutants have not fixed:
$p_{surv}=-\sum_{i=1}^{N}\left[1-(g/f_{R})^{i}\right]\,Ng\mu\,\frac{\pi_{i}}{\pi_{1}}(\tilde{\mathbf{R}}^{-1})_{i1}\,.$ (S16)

### v2

@@ Results
We explore the impact of spatial structure on population survival upon drug addition in three structures with more reduced symmetry than the clique in the Supplementary Appendix Section 4. First, we consider a two-dimensional square lattice with periodic boundary conditions. Second, we consider a star comprising a central deme connected to $D-1$ leaves, where all leaves are equivalent [marrec_toward_2021]. Third, we consider a line (or linear stepping-stone model) composed of $D$ demes, where each deme is connected to two neighbors, except at the two ends [servajean_impact_2024]. In all cases, we choose the migration rates so that the total incoming migration rate to a deme is the same as in the clique, with the exception of the central deme in the star with asymmetric migrations, and of the end demes in the line with asymmetric migrations. This makes the re-invasion of an R deme by S bacteria as similar as possible across structures, but not fully. Our simulation results in the Supplementary Appendix Section 4 show that all these structured populations have the same probability to survive treatment if $t_{\textrm{add}}<\langle t_{af\,F}(N^{*},D)\rangle$, i.e., in cases where usually not more than one deme has fixed resistance before drug addition. For larger values of drug addition time $t_{\textrm{add}}$, our results for the star depend on migration asymmetry and significantly differ from those obtained for other structures, although the difference remains minor. We interpret this difference as arising from S individuals possibly re-invading demes where R mutants have fixed, which remains structure-dependent.
Overall, our main result that spatial structure increases survival probability upon drug addition is qualitatively robust to changing the graph on which demes are placed, as it largely depends on local fixation of R mutants. However, migration patterns impact the re-invasion of R demes by S bacteria, and because of this, the conditions on migration rate we identified in Eqs 5 and 6 for the clique will differ for other graphs. The spread of R mutants to other demes by migration before drug addition, and the colonization of the graph by R mutants after drug is added, will also be affected by graph structure.
Inoculum that contains R mutants.
So far, we focused on a sensitive inoculum and on resistant individuals appearing through mutations (see “Models and methods”). However, resistant mutants can be present in the inoculum. This can be promoted e.g. by the presence of antibiotic in other hosts or in the environment. Furthermore, spatially structured populations starting with a fraction of R mutants in each deme were recently considered in in vitro experiments and modeling in Ref. [kreger_role_2023]. To connect to these cases, we consider the initial condition where some mutants are already present in each deme of a clique, see Supplementary Appendix Section 5. We observe that in this case too, spatial structure increases the survival probability of the population. Indeed, mutant extinction in all demes is less likely than mutant extinction in a well-mixed population.
@@ Discussion
Here, we showed that spatial structure can facilitate the survival of a bacterial population to antibiotic treatment, starting from a sensitive inoculum. Indeed, the bacterial population can be rescued if antibiotic resistant mutants are present when drug is added. While the emergence of resistant bacteria by random mutations only depends on total population size and not on spatial structure, their fate can be affected by spatial structure. If the mutation that provides resistance is neutral or deleterious, which is usually the case, its probability of fixation is increased in smaller populations. This leads to local fixation of resistant mutants in demes, which then constitute refugia of resistant mutants and allow the population to survive when drug is added. Because of this, spatial structure facilitates evolutionary rescue by drug resistance. The survival probability of the population increases when the migration rate between demes is decreased, and when the number of demes is increased. We showed that this effect exists in the rare mutation regime, for relatively rare to rare migration rates, and we quantified these conditions. While it holds for all costs of resistance, the timescales involved become very long for substantial costs. Thus, we mainly focused on neutral resistant mutants and those with moderate costs. After drug is added and the population is rescued by resistance, migrations allow resistant mutants to colonize all demes. While we considered a minimal model where all demes are equivalent and connected to each other with identical migration rates, bacterial populations can have diverse spatial structures. We showed that our key finding that spatial structure facilitates rescue by resistance still holds for more complex spatial structures, and when resistant mutants are present in the inoculum, as it relies on local fixation.
@@
Because local fixation of neutral resistant mutants is instrumental to our effect, it involves rather long timescales. For neutral resistant mutants, the average appearance time $\langle t_{af\,F}\rangle=1/(D\mu g)$ of a locally successful mutant in the fastest deme is of particular relevance. As it scales as the inverse mutation probability per individual and per generation $1/\mu$, it can be quite long, even considering that there are often several different mutational targets that give rise to resistance, leading to a larger effective $\mu$. However, this timescale is inversely proportional to the number $D$ of demes, meaning that it can potentially become arbitrarily small when increasing population subdivision. Recall however that we need to be in the rare mutation regime for our effect to be important. With a mutation probability per nucleotide and per generation of $10^{-10}$ [lee_rate_2012], assuming 10 possible mutations for the development of resistance to a given drug, gives $\mu=10^{-9}$. Then, for bacteria dividing once per hour, we find $\langle t_{af\,F}\rangle=10^{8}$ h for $D=10$, but $\langle t_{af\,F}\rangle=10^{3}$ h for $D=10^{6}$. In the latter case, we would be in the rare mutation regime for deme sizes up to a few hundred bacteria.
[…]
In this work, we focused on a perfect biostatic drug that prevents any division of sensitive bacteria. However, our main findings generalize to biocidal drugs that increase the death rate of sensitive bacteria, or to drugs which combine both modes of actions. Indeed, what happens before the addition of drug, in particular the local fixation of resistant mutants, is not impacted by the mode of action of the drug. Thus, the effect of spatial structure we evidenced here holds independently of this. The difference between the biostatic and the biocidal case lies in the decay of the sensitive bacteria once drug is added. In the biocidal case, sensitive microbes may still divide during this phase, and new resistant mutants may then appear [marrec_resist_2020, czuppon_stochastic_2023]. This is the case both in well-mixed and in structured populations, with only minor differences between them, due to the stochasticity of extinction. Thus, our main findings are robust to the mode of action of the drug. Beyond modes of action, it would be interesting to study the impact of spatial structure on multi-step drug resistance evolution [Igler21].
@@ 1 Derivation of analytical results presented in the main text
We describe population composition using a stochastic model (see “Models and methods” in the main text). This implies that the number $N$ of bacteria in a deme is constantly fluctuating in time. However, because $K\gg 1$ and $g\ll f_{S}=1$, these fluctuations around the steady-state deme size $N^{*}$ remain small. At steady state, the division rate balances the death rate, leading to: $f_{S}(1-N^{*}/K)=g$ in a population that only comprises S bacteria. This yields
$N^{*}=K\left(1-g/f_{S}\right)$ (S1)
In all the analytical calculations that follow, we approximate deme size $N$ by $N^{*}$ (see Ref. [marrec_resist_2020]).
Let us consider a deme where R mutants have not fixed, in the absence of drug. Let us determine the probability $p_{\textrm{pres}}$ of presence of mutants that arise through mutations upon division of sensitive individuals, and whose lineage is destined to go extinct in the absence of drug, but does not undergo rapid stochastic extinction after drug is added. We follow the method exposed in Ref. [marrec_resist_2020], which is based on the transition rate matrix of the process.
Presence of $i$ mutants from a lineage destined to go extinct without drug.
To incorporate the stochastic extinction effect, we need to distinguish each case where exactly $i$ mutants are present, as the probability of stochastic extinction will depend on $i$. The probability of having exactly $i$ mutants present in a well-mixed population with $N^{*}$ individuals reads $p_{R}(i)=\tau_{R}^{d}(i)\,N^{*}\mu g$, where $\tau_{R}^{d}(i)$ is the average time spent in a state with $i$ mutants by a resistant lineage destined for extinction (known as the sojourn time [ewens_mathematical_2004]), and $N^{*}\mu g$ the total mutation rate, with $g$ the death rate and $\mu$ the mutation probability upon division. We have $\tau_{R}^{d}(i)=-\pi_{i}/\pi_{1}(\mathbf{\tilde{R}}^{-1})_{i1}$, where $\pi_{i}$ is the probability that resistant mutants go extinct, starting from $i$ of them, while $\mathbf{\tilde{R}}$ is the reduced transition rate matrix for mutants (i.e., the transition rate matrix where rows and columns corresponding to absorbing states are eliminated), see Ref. [marrec_resist_2020]. This yields the probability that $i$ mutants from a lineage destined to go extinct are present in the deme:
$p_{R}(i)=-N^{*}g\mu\,\frac{\pi_{i}}{\pi_{1}}(\tilde{\mathbf{R}}^{-1})_{i1}\,.$ (S2)
As discussed in the main text, when drug is added to the system, the population can survive if mutants are present when the drug is added. For survival, it is also necessary that the mutant lineages do not quickly go stochastically extinct after drug is added. Let us thus calculate the probability that such a stochastic extinction occurs.
Probability of stochastic extinction.
Let us consider one resistant bacterium with replication rate $f_{R}$ and death rate $g$. Let us denote by $e$ the extinction probability of its lineage. This resistant individual can die without dividing with probability $p_{D}=g/(f_{R}+g)$, giving an extinction probability of $e_{D}=1$. It can also divide before dying with probability $p_{R}=f_{R}/(f_{R}+g)$, leading to the extinction probability $e_{R}=e^{2}$, assuming independence of lineages, in the branching process framework [harris_branching_1963]. We can then write an equation for the extinction probability $e$ of the lineage of the single resistant bacterium:
$e=p_{D}\,e_{D}+p_{R}\,e_{R}=\frac{g}{f_{R}+g}+e^{2}\frac{f_{R}}{f_{R}+g}.$ (S3)
The solutions to this equation are $e=1$ or $e=g/f_{R}$ if $g<f_{R}$. Thus, assuming again that the fate of each lineage is independent, we obtain an extinction probability $(g/f_{R})^{i}$ starting from $i$ mutants.
Presence of a mutant lineage destined to go extinct without drug, but that does not undergo stochastic extinction when drug is added.
We finally combine Eq S2 with the stochastic extinction probability to find:
$p_{\textrm{pres}}=\sum_{i=1}^{N^{*}}\left[1-(g/f_{R})^{i}\right]p_{R}(i)=-\sum_{i=1}^{N^{*}}\left[1-(g/f_{R})^{i}\right]\,N^{*}g\mu\,\frac{\pi_{i}}{\pi_{1}}(\tilde{\mathbf{R}}^{-1})_{i1}\,.$ (S4)
Eq S4 is a good approximation of the survival probability of the population on timescales smaller than the average time of appearance of a successful R mutant, since R mutants that appear before are doomed to go extinct in the absence of drug.
Lifetime of a mutant lineage destined to go extinct without drug.
In the absence of drug, the total average lifetime $\tau_{R}^{d}$ of a mutant lineage destined to go extinct can be obtained by summing the sojourn times $\tau_{R}^{d}(i)$ expressed above (see also Ref. [ewens_mathematical_2004]):
$\tau_{R}^{d}=\sum_{i=1}^{N^{*}}\tau_{R}^{d}(i)=-\sum_{i=1}^{N^{*}}\frac{\pi_{i}}{\pi_{1}}(\mathbf{\tilde{R}}^{-1})_{i1}\,.$ (S5)
For neutral R mutants, this gives
$\tau_{R}^{d}=\frac{1}{g}\left(\frac{N^{*}}{N^{*}-1}\sum_{i=1}^{N^{*}-1}\frac{1}{i}-1\right)\approx\frac{1}{g}\log(N^{*})\,,$ (S6)
where we assumed $N^{*}\gg 1$ to obtain the last expression. This lifetime is only weakly impacted by population size.
@@ 1.2 Appearance of a mutant that fixes in a well-mixed population
In our model, mutations from sensitive to resistant individuals happen at birth with probability $\mu$. Since at equilibrium $f_{S}\left(1-N^{*}/K\right)=g$, the average time it takes for a mutant to appear in a population of $N^{*}$ wild-types is in the rare mutation regime:
$\langle t_{\textrm{app}}(N^{*})\rangle=\frac{1}{N^{*}\mu g}.$ (S7)
Thus, the average time $\langle t_{af\,W}(N^{*})\rangle$ of appearance of a mutant that fixes in a well-mixed population of fixed size $N^{*}$ is given by:
$\langle t_{af\,W}(N^{*})\rangle=\langle t_{\textrm{app}}(N^{*})\rangle\times\frac{1}{p_{\textrm{fix}}(N^{*})}=\frac{1}{N^{*}\mu g\,p_{\textrm{fix}}(N^{*})}.$ (S8)
Cost-free mutants.
The fixation probability of neutral mutants in a population of size $N^{*}$ is:
$p_{\textrm{fix}}(N^{*})=\frac{1}{N^{*}}\,.$ (S9)
Thus, the average time $\langle t_{af\,W}\rangle$ of appearance of a neutral mutant that fixes in a well-mixed population of fixed size $N^{*}$ is given by:
$\langle t_{af\,W}(N^{*})\rangle=\langle t_{\textrm{app}}(N^{*})\rangle\times\frac{1}{p_{\textrm{fix}}(N^{*})}=\frac{1}{\mu g}\,.$ (S10)
This implies that for cost-free mutants, the average time of appearance of a successful mutant is independent of the population size: it remains the same for a single deme and a well-mixed population.
Mutants with cost $\delta$.
More generally, the fixation probability of a mutant with fitness cost $\delta$ in a well-mixed population of fixed size $N^{*}$ described by the Moran process is [ewens_mathematical_2004]
$p_{\textrm{fix}}(N^{*})=\frac{(1-\delta)^{-1}-1}{(1-\delta)^{-N^{*}}-1}\,.$ (S11)
If $\delta\ll 1/N^{*}$, then to leading order
$p_{\textrm{fix}}(N^{*})=\frac{1}{N^{*}}\,.$ (S12)
The mutant is then said to be effectively neutral. Conversely, if $\delta\gg 1/N^{*}$, then to leading order
$p_{\textrm{fix}}(N^{*})=\delta e^{-N^{*}\delta}\,.$ (S13)
Mutants with such a substantial fitness cost have a fixation probability that is exponentially suppressed.
In Fig 2B, as $K\delta=1$, we employ the most general formula for $p_{\textrm{fix}}$, see Eq S11. The cost regimes leading to the simplified expressions of $p_{\textrm{fix}}$ in Eqs S12 and S13 are further discussed in Section 3.2.
@@ 1.3 Appearance of the slowest mutant that fixes in a structured population
Let us now focus on the appearance of a mutant that fixes in the slowest deme in a population composed of $D$ independent demes which each have steady-state size $N^{*}$, in the absence of migrations. By slowest deme, we mean the deme in which a locally successful mutant (i.e. a mutant destined to fix locally in its deme of origin) takes the longest time to appear.
For a mutant to fix in the slowest deme, mutants must have fixed in all others. Let us denote by $t_{af\,S}$ (resp. $t_{af\,d}$) the time of appearance of a mutant destined to fix in the slowest deme (resp. in any deme). The probability $P(t_{af\,S}\leq t)$ that $t_{af\,S}$ is smaller or equal than $t$ reads for any $t$:
@@
The appearance of a locally successful mutant is a Poisson process with rate $\lambda=N^{*}\mu g\,p_{\textrm{fix}}(N^{*})$ (corresponding to the inverse of Eq S8). Thus, the time $t_{af\,d}$ is exponentially distributed with rate $\lambda$ and probability density
$p_{d}(t_{af\,d}=t)=\lambda e^{-\lambda t}\,,$ (S15)
and we have $P(t_{af\,d}\leq t)=1-e^{-\lambda t}$. This allows us to express the probability density $p_{S}(t)$ of appearance of a mutant in the slowest deme:
$\begin{split}p_{S}(t)dt&=dp(t_{af\,S}\in[t,t+dt])=P(t_{af\,S}\leq t+dt)-P(t_{af\,S}\leq t)\\ &=\frac{dP(t_{af\,S}\leq t)}{dt}dt=D\left[P(t_{af\,d}\leq t)\right]^{D-1}\,\frac{dP(t_{af\,d}\leq t)}{dt}dt\\ &=D\left[1-e^{-\lambda t}\right]^{D-1}\lambda e^{-\lambda t}dt\,.\end{split}$ (S16)
@@ 1.4 Appearance of the fastest mutant that fixes in a structured population
$\begin{split}p_{F}(t)dt&=dp(t_{af\,F}\in[t,t+dt])=-\frac{dP(t_{af\,F}>t)}{dt}dt\\ &=-D[P(t_{af\,d}>t)]^{D-1}\frac{dP(t_{af\,d}>t)}{dt}dt\\ &=-D[1-P(t_{af\,d}\leq t)]^{D-1}\left[-\frac{dP(t_{af\,d}\leq t)}{dt}\right]dt\\ &=D\lambda e^{-\lambda Dt}dt\,,\end{split}$ (S19)
with $\lambda=N^{*}\mu g\,p_{\textrm{fix}}(N^{*})$, as previously defined. Note that $p_{F}(t)$ is positive for all $t$ and normalized, as expected.
@@
In the case of cost-free mutants, using Eq S9 gives
$\langle t_{af\,F}(N^{*},D)\rangle=\frac{1}{D\mu g}\,,$ (S21)
In the case where resistance carries a cost, using Eq S11 gives
$\langle t_{af\,F}(N^{*},D)\rangle=\frac{1}{DN^{*}\mu g}\frac{(1-\delta)^{-N^{*}}-1}{(1-\delta)^{-1}-1}\,.$ (S22)
@@ 1.5 Survival probability of the population with $\gamma=0$
We checked that Eq S23 was satisfied in our numerical simulations, within the errors given by the standard error of the proportion.
1.6 Closed form of the survival probability for well-mixed and fully subdivided populations
In our structured populations, mutants appear at random in any deme. The population can survive in two distinct ways:
  • Due to the presence of successful lineages that have fixed locally or globally in the population, and allow survival when drug is added.
  • Due to the presence of a mutant lineage destined for extinction in the absence of drug at the time $t_{\textrm{add}}$ when drug is added. This mutant lineage can proliferate in the presence of drug. The population then survives, unless this mutant lineage undergoes a quick stochastic extinction.
The survival probability can be then written by summing over these two distinct scenarios:
$\begin{split}p_{\textrm{surv}}(t_{\textrm{add}})&=p_{\textrm{succ, surv}}(t_{\textrm{add}})+p_{\textrm{no\, succ, surv}}(t_{\textrm{add}})\\ &=p_{\textrm{succ}}(t_{\textrm{add}})\,p_{\textrm{surv}|\textrm{succ}}+\left[1-p_{\textrm{succ}}(t_{\textrm{add}})\right]p_{\textrm{surv}|\textrm{no\, succ}}\\ &=p_{\textrm{succ}}(t_{\textrm{add}})+\left[1-p_{\textrm{succ}}(t_{\textrm{add}})\right]p_{\textrm{pres}}\,.\end{split}$ (S26)
In the first line, we denoted by $p_{\textrm{succ, surv}}(t_{\textrm{add}})$ the probability that a successful mutant has appeared by $t_{\textrm{add}}$ and leads to population survival, and by $p_{\textrm{no\, succ, surv}}(t_{\textrm{add}})$ the probability that no successful mutant has appeared at $t_{\textrm{add}}$ but that the population survives (thanks to the presence of a mutant lineage that was destined for extinction). In the second line, $p_{\textrm{succ}}(t_{\textrm{add}})$ is the probability that a successful mutant has appeared by $t_{\textrm{add}}$. The probability $p_{\textrm{surv}|\textrm{succ}}$ of survival conditioned on the presence of a successful mutant is one, while the probability $p_{\textrm{surv}|\textrm{no\, succ}}$ of survival conditioned on the absence of any successful mutant is $p_{\textrm{pres}}$, given by Eq S4.
Let us now calculate $p_{\textrm{succ}}(t_{\textrm{add}})$ in a well-mixed population of fixed size $N^{*}$. Given that the appearance of a successful mutant is a Poisson process with rate $\lambda=N^{*}\mu gp_{\textrm{fix}}(N^{*})$, the associated time of appearance is distributed according to Eq S15. Thus, $p_{\textrm{succ}}(t_{\textrm{add}})$ reads:
[…]
For neutral mutants in a well-mixed population of size $N^{*}$, this reads:
$p_{\textrm{succ}}(t_{\textrm{add}})=1-e^{-\mu gt_{\textrm{add}}},$ (S28)
while for deleterious mutants in a well-mixed population of size $N^{*}$ we have:
$p_{\textrm{succ}}(t_{\textrm{add}})=1-\exp\left(-N^{*}\mu g\frac{(1-\delta)^{-1}-1}{(1-\delta)^{-N*}-1}t_{\textrm{add}}\right).$ (S29)
Let us now consider a structured population, and ask whether a locally successful mutant has appeared by $t_{\textrm{add}}$. For the appearance time of successful mutants in the fastest deme of a structured population, the probability density is given by Eq S19, yielding:
[…]
If mutants are neutral, we have:
$p_{\textrm{succ}}(t_{\textrm{add}})=1-e^{-D\mu gt_{\textrm{add}}},$ (S31)
and if they carry a resistance cost:
$p_{\textrm{succ}}(t_{\textrm{add}})=1-\exp\left(-DN^{*}\mu g\frac{(1-\delta)^{-1}-1}{(1-\delta)^{-N*}-1}t_{\textrm{add}}\right).$ (S32)